OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 3/2x + 4 and passes through the point (−6, 3).

OpenStudy (anonymous):

y = (-2/3)x + b 3 = (-2/3)(-6) + b b = -1 The equation of the line perpendicular to the line y= (3/2)x+4 is y=(-2/3)x - 1 I think one second I am checking my answer

OpenStudy (anonymous):

Yeah that should be right. Do you see how I did this?

OpenStudy (anonymous):

yeah thanks man.

OpenStudy (anonymous):

what about this next one?

OpenStudy (anonymous):

What next one?